A car travels at 50 mph for 2 hours. It then travels another 200 miles at an average of 60 mph. What was the car's average speed for the entire trip?
1
Expert's answer
2020-09-01T11:01:39-0400

"Distance= speed\times time". For the first part of the journey distance will be "d=50mph\times 2hrs= 100miles."


Total time taken for second part of the journey="\frac{200}{60}=3.33hrs."


We know that ="Average speed=\frac{total distance}{total time taken}". Therefore,


"Total distance= 100miles+200miles= 300miles."


"Total time taken= 2hrs+3.3hrs=5.3hrs."


Thus,"Average speed= \frac{300}{5.3}= 56.6mph"